# Cosmic (super)strings with a time-varying tension

### v1

@@ 4.2.2 Dynamical system analysis
The evolution equations (56)-(57) can be recast as the autonomous system
$\begin{cases}\frac{{\rm d}\xi}{{\rm d}z}=\xi\left[\left(\frac{2}{n}-1\right)+\frac{v^{2}}{2}\left(\frac{4}{n}-m\right)\right]+\frac{\tilde{c}(v)v}{2}\\ \frac{{\rm d}v}{{\rm d}z}=(1-v^{2})\left[\frac{k(v)}{\xi}-\left(\frac{4}{n}-m\right)v\right],\\ \end{cases}$ (59)
through the change of variable $z=\log t$ and $L=\xi(t)t$. As in the previous sections, we assume a universe dominated by a background fluid with $\rho\sim a^{-n}$ and $\mu\sim t^{-m}$, with $n>2$ and $m>0$. Note that in the first equation we have not included the gravitational back-reaction, i.e. the effect of GW production, which gave a sub-leading effect in the contracting universe analysis of Avelino:2002hx ; Avelino:2002xy . In our case, due to the decreasing tension, we expect that the effect will be negligible. Moreover, the attentive reader will have noticed that the chopping parameter $\tilde{c}$ has been promoted to a velocity-dependent function $\tilde{c}(v)$. Although so far we have assumed $\tilde{c}$ to be a constant, for $mn>4$ the network is expected to probe the ultra-relativistic regime, where relativistic corrections become important. Taking Lorentz contraction of the correlation length into account, a reasonable ansatz is $\tilde{c}(v)\simeq\tilde{c}(1-v^{2})^{1/2}$ Avelino:2002hx ; Avelino:2002xy , which implies $\tilde{c}\rightarrow 0$ in the ultra-relativistic limit. However, there may also be additional loss terms in the first equation (for example decays to other particles) that we have not taken into account, potentially leading to a small, effective $\tilde{c}\neq 0$. For this reason, following Avelino:2002hx ; Avelino:2002xy , we will also consider the case where $\tilde{c}(v)$ asymptotes to a constant $\tilde{c}_{\infty}\neq 0$ as $v\rightarrow 1$.
In this form, (59) can be analysed with standard techniques for dynamical systems. In particular, one can distinguish two different classes of fixed points, examined in detailed below. The first one is that of scaling fixed points, which straightforwardly generalise the scaling solution for $m=0$. The second class consists of ultra-relativistic fixed points, where $v$ asymptotes to $1$. These are reminiscent of what happens for a network of cosmic strings in a contracting universe Avelino:2002hx ; Avelino:2002xy , with a few differences that will be reviewed. Here we will just state the main outcome of the analysis, leaving the detailed proofs to Appendix C.
@@
$\xi=\sqrt{\frac{n}{8}\frac{k(\bar{v})\left(k(\bar{v})+\tilde{c}\right)}{\left(1-\frac{mn}{4}\right)\left(1-\frac{2}{n}\right)}}\quad\quad\quad\quad\bar{v}=\sqrt{\frac{n}{2}\frac{k(\bar{v})}{k(\bar{v})+\tilde{c}}\frac{1-\frac{2}{n}}{1-\frac{mn}{4}}},$ (60)
which reduce to (30) and (31) for $m=0$. The multiplicity of the solutions (60) depends on the possible values of the velocity $\bar{v}$, determined by the solutions to
$k(\bar{v})-\frac{\tilde{c}\bar{v}^{2}}{\frac{2(n-2)}{4-mn}-\bar{v}^{2}}=0.$ (61)
For $mn<4$ there is always one such fixed point, which has $\bar{v}^{2}<\frac{1}{2}$ and is always an attractor. It corresponds to the usual fixed point found for $m=0$. For $mn>4$, solutions to (61) always come in pairs for a momentum parameter of the form (33). However, only the one with lowest value of $v^{2}$ will be stable, while the other one is always a saddle. From (60), it is also easy to see that if $mn>4$, such solutions must have $k(\bar{v})\leq 0$, while $k(\bar{v})\geq 0$ if $mn<4$. As we have just mentioned, this traces back to the fact that, if $02H+\frac{1}{\mu}\frac{{\rm d}\mu}{{\rm d}t}<$ 0, the RHS of (57) will always be positive, and $v$ will keep growing until $k(v)$ becomes negative. [^17] Notice that this is the same condition required for loops to grow faster than the scale factor, as shown in Conlon:2024uob . In particular, the phenomenological forms of $k(v)$ proposed in the literature (See Eq.(33)), which have been validated through numerical simulations, are all strictly positive for $v^{2}\leq\frac{1}{2}$. In section 4.1.2, we also provided analytic evidence that small scale structure induces a negative $k(v)$ whenever $mn>4$.
@@
$v_{r}=1\quad\quad\quad\xi_{r}=\frac{\tilde{c}_{\infty}n}{2(n-2)n+mn-4}.$ (62)
@@
$mn>{\rm Max}\left\{4,4+n\left(\frac{m}{2}-1\right)\right\},$ (63)
so it can never be an attractor for $mn<4$. The solution to (62) is of the genuine scaling type, so its energy density will behave as in (58). However, if $\tilde{c}\ll\xi_{0}$, as expected in practical applications, trajectories approaching the fixed point will effectively behave as if $\tilde{c}_{\infty}=0$ for a time
$\frac{t_{r}}{t_{0}}\lesssim\left(\frac{\xi_{0}}{\tilde{c}_{\infty}}\frac{2(n-2)+mn-4}{n}\right)^{\frac{2n}{2(n-2)+mn-4}}.$ (64)
In this timeframe, the network behaves like a gas of very relativistic strings, with an equation of state equal to that of radiation and $\rho\sim a^{-4}$. This is analogous to the case of cosmic strings in a contracting universe Avelino:2002hx ; Avelino:2002xy . However, a crucial difference is that in this case it can coexist with a stable scaling fixed point, and the outcome of the system crucially depends on the initial conditions, as well as the values of the various parameters.
@@
If $\tilde{c}(v)$ approaches 00 in the ultra-relativistic regime, the results of the previous paragraph cannot be applied, as $\xi=0$ is a singular point of the system (59). However, for a chopping parameter of the form $\tilde{c}(1-v^{2})^{\alpha}$, it can be shown that there still exists a fixed point with $v=1$ and $\xi\rightarrow 0$. For $\alpha=\frac{1}{2}$, as motivated above, the latter is an attractor if
$m>\frac{16+2n}{5n}.$ (65)
@@
$\xi(t)=\xi_{0}\left(\frac{t}{t_{0}}\right)^{-\frac{1}{2n}\left[2(n-2)+mn-4\right]}\quad\quad\text{and}\quad\quad\rho=\frac{\mu(t)}{L(t)^{2}}\sim\frac{1}{t^{8/n}}\sim\frac{1}{a(t)^{4}},$ (66)
@@ 4.2.3 Dependence on initial conditions
Therefore, the evolution of the system will be heavily dependent on initial conditions, and the timescales one is interested in. In this last section, we would like to give a few examples for realistic values of the parameters. For the rest of the section, we will further assume $\tilde{c}(v)=\tilde{c}(1-v^{2})^{1/2}$, with $\tilde{c}=0.23$ Martins:2000cs . We can start from the phenomenologically interesting case $m=1,n=6$, which corresponds to volume kination in a string-theoretic embedding. In Figures 1(a)-1(b), we plotted regions where the late-time value of $v$ is close to either critical points after a time $t=e^{10}$ and $t=e^{20}$, [^19] By convention, we have arbitrarily defined this as $|v-\bar{v}|<0.01$. as a function of initial conditions for $\xi$ and $v$. The difference between the two plots shows how it may take a relatively long time to reach one of the attractors, and even how trajectories that stay very close to one of the fixed point for a long time can still end up converging to the other one. At network formation, the density of strings per Hubble patch is of order $\mathcal{O}(1)$, and we expect $\xi\sim\mathcal{O}(1)$. Therefore, the proper scaling fixed point will most likely be reached at very late times, unless the initial velocity is very close to $1$.
@@ 4.3 Loop evolution
Having concluded that a scaling regime is achieved when the tension varies, we now study the averaged dynamics of subhorizon loops. For a single loop, one can similarly integrate (42) to obtain the evolution of its total energy $E_{\ell}\equiv\mu\ell$:
$\frac{{\rm d}E_{\ell}}{{\rm d}t}=E_{\ell}H(1-2v_{\ell}^{2})+E_{\ell}\frac{1}{\mu}\frac{d\mu}{dt}(1-v^{2})-\Gamma^{\prime}G\mu^{2}v_{\ell}^{6},$ (67)
@@
$\frac{{\rm d}E_{\ell}}{{\rm d}t}=E_{l}\frac{1}{2\mu}\frac{d\mu}{dt}-\Gamma G\mu^{2},$ (68)
which matches with Eq. (34) of Emond:2021vts for $\mu\sim\frac{1}{a^{2}}$. For a decreasing tension, the energy of the loops will also decrease with time as well. It also follows that the time evolution of the loop length is given by
$\frac{{\rm d}\ell}{{\rm d}t}=-\ell\frac{1}{2\mu}\frac{d\mu}{dt}-\Gamma G\mu,$ (69)
where one of the terms has (crucially) changed sign, now allowing for the loop to grow even if it is emitting gravitational radiation. Since $\ell\lesssim t^{m/2}$, for $m<2$ the loops will always grow slower than the horizon size. This justifies a posteriori the assumptions that $v_{\ell}^{2}=\frac{1}{2}$, true on sub-horizon scales. Notice that if the tension is decreasing with time, the first contribution to (69) is positive, and if it dominates over energy loss loops can in principle grow with time. With the specific dependence $\mu(t)=\mu_{0}\left(\frac{t_{0}}{t}\right)^{m}$, this is solved by
$\ell(t)=\left(\frac{t}{t_{i}}\right)^{m/2}\left(\alpha-\frac{\Gamma G\mu_{0}}{\frac{3m}{2}-1}\left(\frac{t_{0}}{t_{i}}\right)^{m}\right)t_{i}+\frac{\Gamma G\mu_{0}}{\frac{3m}{2}-1}\left(\frac{t_{0}}{t}\right)^{m-1}t_{0}\,,$ (70)
@@
$\ell(t)=\left[\alpha-\Gamma G\mu_{0}\left(\frac{t_{0}}{t_{i}}\right)^{2/3}\log\left(\frac{t}{t_{i}}\right)\right]\left(\frac{t}{t_{i}}\right)^{1/3}t_{i}\,.$ (71)
@@ 4.4 Percolation
As a consequence of the decreasing tension, it was noticed in the recent paper Conlon:2024uob how loops on sub-horizon lengths will grow in size more rapidly than the scale factor if $02\mathcal{H}+\frac{\dot{\mu}}{\mu}<$ 0 is satisfied. If their density is large enough, such loops can merge with each other to form larger ones and eventually percolate, giving rise to a network of long strings. While such a mechanism was introduced in Conlon:2024uob to provide an alternative scenario for the formation of a cosmic (super)string network, one might ask if percolation itself may not make a pre-existing string network unstable. Indeed, as we have seen in the previous sections the equilibrium (scaling) configuration of any cosmic string network contains a population of loops arising from the reconnection of long strings. If such loops were able to percolate, and convert their energy back into a network of long strings, one might be worried that they would alter the properties of the network, and perhaps even overclose the universe.
For this reason, let us estimate how long it would take for percolation to occur, assuming the existence of a network from the time $t_{0}$. For a fixed final time $t$, the length distribution of the loop population can be derived by the production rate (38), as well as the fact that a loop produced at the time $t_{p}$ will have grown as
$\ell(t)=\alpha t_{p}\left(\frac{t}{t_{p}}\right)^{m/2},$ (72)
@@
$\frac{{\rm d}n}{{\rm d}\ell}=\frac{2\mathcal{F}_{\alpha}\alpha^{\frac{6}{2-m}}}{(2-m)}\frac{c\bar{v}}{\gamma\xi^{3}}\frac{t^{\frac{3m}{2-m}}}{\ell^{\frac{8-m}{2-m}}}.$ (73)
As in Conlon:2024uob , percolation may be estimated to have occured when the typical string separation is the same order as the string length, namely
$\int_{\ell(t_{0})}^{\ell(t)}{{\rm d}\ell}\,\frac{\ell(t)^{3}}{a(t)^{3}}\frac{{\rm d}n}{{\rm d}\ell}\sim\mathcal{O}(1).$ (74)
@@
$\frac{t}{t_{0}}\lesssim\left[\frac{3m}{2\mathcal{F}_{\alpha}}\frac{\gamma\xi^{3}}{c\bar{v}}\alpha^{\frac{3m-6}{2-m}}\right]^{\frac{2n}{3(mn-4)}}.$ (75)
@@
$\frac{t}{t_{0}}\lesssim 10^{6}-10^{7}.$ (76)
This is a sizeable fraction (on a log scale) of the maximal allowed lenght of kination in LVS, $\frac{t}{t_{0}}\sim 10^{10}$ Conlon:2024uob .
The question of what happens afterwards is slightly more subtle, and we suspect it cannot be answered with the tools at our disposal. [^20] Progress on this matter could be achieved using the Boltzmann equation approach of Copeland:1998na ; Frey:2023khe . Therefore, this paragraph should be considered mainly as speculative. As the loops merge with one another and start to percolate, they start to feed back into the population of long strings. To model this within the VOS framework, a cheap attempt may be to decrease the loop chopping efficiency $\tilde{c}$ governing the decays of long strings into loops. If $\tilde{c}$ remains lower bounded, then the network will also be exhibiting a quasi-scaling regime as per Eq. (125). On the other hand, $L(t)$ may grow sub-linearly if $\tilde{c}\rightarrow 0$. Even if this is the case, the network can never dominate the energy density of the universe, unlike what would happen with a constant tension. Indeed, by (125),
$\rho=\frac{\mu}{L^{2}}\leq\frac{C}{t^{2}},$ (77)

### v2

@@ 4.2.2 Dynamical system analysis
The evolution equations (4.18)-(4.19) can be recast as the autonomous system
$\begin{cases}\frac{{\rm d}\xi}{{\rm d}z}=\xi\left[\left(\frac{2}{n}-1\right)+\frac{v^{2}}{2}\left(\frac{4}{n}-m\right)\right]+\frac{\tilde{c}(v)v}{2}\\ \frac{{\rm d}v}{{\rm d}z}=(1-v^{2})\left[\frac{k(v)}{\xi}-\left(\frac{4}{n}-m\right)v\right],\\ \end{cases}$ (4.21)
through the change of variable $z=\log t$ and $L=\xi(t)t$. As in the previous sections, we assume a universe dominated by a background fluid with $\rho\sim a^{-n}$ and $\mu\sim t^{-m}$, with $n>2$ and $m>0$. Note that in the first equation we have not included the gravitational back-reaction, i.e. the effect of GW production, which gave a sub-leading effect in the contracting universe analysis of [98, 99]. In our case, due to the decreasing tension, we expect that the effect will be negligible. Moreover, the attentive reader will have noticed that the chopping parameter $\tilde{c}$ has been promoted to a velocity-dependent function $\tilde{c}(v)$. Although so far we have assumed $\tilde{c}$ to be a constant, for $mn>4$ the network is expected to probe the ultra-relativistic regime, where relativistic corrections become important. Taking Lorentz contraction of the correlation length into account, a reasonable ansatz is $\tilde{c}(v)\simeq\tilde{c}(1-v^{2})^{1/2}$ [98, 99], which implies $\tilde{c}\rightarrow 0$ in the ultra-relativistic limit. However, there may also be additional loss terms in the first equation (for example decays to other particles) that we have not taken into account, potentially leading to a small, effective $\tilde{c}\neq 0$. For this reason, following [98, 99], we will also consider the case where $\tilde{c}(v)$ asymptotes to a constant $\tilde{c}_{\infty}\neq 0$ as $v\rightarrow 1$.
In this form, (4.21) can be analysed with standard techniques for dynamical systems. In particular, one can distinguish two different classes of fixed points, examined in detailed below. The first one is that of scaling fixed points, which straightforwardly generalise the scaling solution for $m=0$. The second class consists of ultra-relativistic fixed points, where $v$ asymptotes to $1$. These are reminiscent of what happens for a network of cosmic strings in a contracting universe [98, 99], with a few differences that will be reviewed. Here we will just state the main outcome of the analysis, leaving the detailed proofs to Appendix C.
@@
$\xi=\sqrt{\frac{n}{8}\frac{k(\bar{v})\left(k(\bar{v})+\tilde{c}\right)}{\left(1-\frac{mn}{4}\right)\left(1-\frac{2}{n}\right)}}\quad\quad\quad\quad\bar{v}=\sqrt{\frac{n}{2}\frac{k(\bar{v})}{k(\bar{v})+\tilde{c}}\frac{1-\frac{2}{n}}{1-\frac{mn}{4}}},$ (4.22)
which reduce to (3.11) and (3.12) for $m=0$. The multiplicity of the solutions (4.22) depends on the possible values of the velocity $\bar{v}$, determined by the solutions to
$k(\bar{v})-\frac{\tilde{c}\bar{v}^{2}}{\frac{2(n-2)}{4-mn}-\bar{v}^{2}}=0.$ (4.23)
For $mn<4$ there is always one such fixed point, which has $\bar{v}^{2}<\frac{1}{2}$ and is always an attractor. It corresponds to the usual fixed point found for $m=0$. For $mn>4$, solutions to (4.23) always come in pairs for a momentum parameter of the form (3.14). However, only the one with lowest value of $v^{2}$ will be stable, while the other one is always a saddle. From (4.22), it is also easy to see that if $mn>4$, such solutions must have $k(\bar{v})\leq 0$, while $k(\bar{v})\geq 0$ if $mn<4$. As we have just mentioned, this traces back to the fact that, if $02H+\frac{1}{\mu}\frac{{\rm d}\mu}{{\rm d}t}<$ 0, the RHS of (4.19) will always be positive, and $v$ will keep growing until $k(v)$ becomes negative. [^19] Notice that this is the same condition required for loops to grow faster than the scale factor, as shown in [35]. In particular, the phenomenological forms of $k(v)$ proposed in the literature (See Eq.(3.14)), which have been validated through numerical simulations, are all strictly positive for $v^{2}\leq\frac{1}{2}$. In section 4.1.2, we also provided analytic evidence that small scale structure induces a negative $k(v)$ whenever $mn>4$.
@@
$v_{r}=1\quad\quad\quad\xi_{r}=\frac{\tilde{c}_{\infty}n}{2(n-2)n+mn-4}.$ (4.24)
@@
$mn>{\rm Max}\left\{4,4+n\left(\frac{m}{2}-1\right)\right\},$ (4.25)
so it can never be an attractor for $mn<4$. The solution to (4.24) is of the genuine scaling type, so its energy density will behave as in (4.20). However, if $\tilde{c}\ll\xi_{0}$, as expected in practical applications, trajectories approaching the fixed point will effectively behave as if $\tilde{c}_{\infty}=0$ for a time
$\frac{t_{r}}{t_{0}}\lesssim\left(\frac{\xi_{0}}{\tilde{c}_{\infty}}\frac{2(n-2)+mn-4}{n}\right)^{\frac{2n}{2(n-2)+mn-4}}.$ (4.26)
In this timeframe, the network behaves like a gas of very relativistic strings, with an equation of state equal to that of radiation and $\rho\sim a^{-4}$. This is analogous to the case of cosmic strings in a contracting universe [98, 99]. However, a crucial difference is that in this case it can coexist with a stable scaling fixed point, and the outcome of the system crucially depends on the initial conditions, as well as the values of the various parameters.
@@
If $\tilde{c}(v)$ approaches 00 in the ultra-relativistic regime, the results of the previous paragraph cannot be applied, as $\xi=0$ is a singular point of the system (4.21). However, for a chopping parameter of the form $\tilde{c}(1-v^{2})^{\alpha}$, it can be shown that there still exists a fixed point with $v=1$ and $\xi\rightarrow 0$. For $\alpha=\frac{1}{2}$, as motivated above, the latter is an attractor if
$m>\frac{16+2n}{5n}.$ (4.27)
@@
$\xi(t)=\xi_{0}\left(\frac{t}{t_{0}}\right)^{-\frac{1}{2n}\left[2(n-2)+mn-4\right]}\quad\quad\text{and}\quad\quad\rho=\frac{\mu(t)}{L(t)^{2}}\sim\frac{1}{t^{8/n}}\sim\frac{1}{a(t)^{4}},$ (4.28)
@@ 4.2.3 Dependence on initial conditions
Therefore, the evolution of the system will be heavily dependent on initial conditions, and the timescales one is interested in. In this last section, we would like to give a few examples for realistic values of the parameters. For the rest of the section, we will further assume $\tilde{c}(v)=\tilde{c}(1-v^{2})^{1/2}$, with $\tilde{c}=0.23$ [37]. We can start from the phenomenologically interesting case $m=1,n=6$, which corresponds to volume kination in a string-theoretic embedding. In Figures 1(a)-1(b), we plotted regions where the late-time value of $v$ is close to either critical points after a time $t=e^{10}$ and $t=e^{20}$, [^21] By convention, we have arbitrarily defined this as $|v-\bar{v}|<0.01$. as a function of initial conditions for $\xi$ and $v$. The difference between the two plots shows how it may take a relatively long time to reach one of the attractors, and even how trajectories that stay very close to one of the fixed point for a long time can still end up converging to the other one. At network formation, the density of strings per Hubble patch is of order $\mathcal{O}(1)$, and we expect $\xi\sim\mathcal{O}(1)$. Therefore, the proper scaling fixed point will most likely be reached at very late times, unless the initial velocity is very close to $1$.
@@ 4.3 Loop evolution
Having concluded that a scaling regime is achieved when the tension varies, we now study the averaged dynamics of subhorizon loops. For a single loop, one can similarly integrate (4.4) to obtain the evolution of its total energy $E_{\ell}\equiv\mu\ell$:
$\frac{{\rm d}E_{\ell}}{{\rm d}t}=E_{\ell}H(1-2v_{\ell}^{2})+E_{\ell}\frac{1}{\mu}\frac{d\mu}{dt}(1-v^{2})-\Gamma^{\prime}G\mu^{2}v_{\ell}^{6},$ (4.29)
@@
$\frac{{\rm d}E_{\ell}}{{\rm d}t}=E_{l}\frac{1}{2\mu}\frac{d\mu}{dt}-\Gamma G\mu^{2},$ (4.30)
which matches with Eq. (34) of [34] for $\mu\sim\frac{1}{a^{2}}$. For a decreasing tension, the energy of the loops will also decrease with time as well. It also follows that the time evolution of the loop length is given by
$\frac{{\rm d}\ell}{{\rm d}t}=-\ell\frac{1}{2\mu}\frac{d\mu}{dt}-\Gamma G\mu,$ (4.31)
where one of the terms has (crucially) changed sign, now allowing for the loop to grow even if it is emitting gravitational radiation. Since $\ell\lesssim t^{m/2}$, for $m<2$ the loops will always grow slower than the horizon size. This justifies a posteriori the assumptions that $v_{\ell}^{2}=\frac{1}{2}$, true on sub-horizon scales. Notice that if the tension is decreasing with time, the first contribution to (4.31) is positive, and if it dominates over energy loss loops can in principle grow with time. With the specific dependence $\mu(t)=\mu_{0}\left(\frac{t_{0}}{t}\right)^{m}$, this is solved by
$\ell(t)=\left(\frac{t}{t_{i}}\right)^{m/2}\left(\alpha-\frac{\Gamma G\mu_{0}}{\frac{3m}{2}-1}\left(\frac{t_{0}}{t_{i}}\right)^{m}\right)t_{i}+\frac{\Gamma G\mu_{0}}{\frac{3m}{2}-1}\left(\frac{t_{0}}{t}\right)^{m-1}t_{0}\,,$ (4.32)
@@
$\ell(t)=\left[\alpha-\Gamma G\mu_{0}\left(\frac{t_{0}}{t_{i}}\right)^{2/3}\log\left(\frac{t}{t_{i}}\right)\right]\left(\frac{t}{t_{i}}\right)^{1/3}t_{i}\,.$ (4.33)
@@ 4.4 Percolation
As a consequence of the decreasing tension, it was noticed in the recent paper [35] how loops on sub-horizon lengths will grow in size more rapidly than the scale factor if $02\mathcal{H}+\frac{\dot{\mu}}{\mu}<$ 0 is satisfied. If their density is large enough, such loops can merge with each other to form larger ones and eventually percolate, giving rise to a network of long strings. While such a mechanism was introduced in [35] to provide an alternative scenario for the formation of a cosmic (super)string network, one might ask if percolation itself may not make a pre-existing string network unstable. Indeed, as we have seen in the previous sections the equilibrium (scaling) configuration of any cosmic string network contains a population of loops arising from the reconnection of long strings. If such loops were able to percolate, and convert their energy back into a network of long strings, one might be worried that they would alter the properties of the network, and perhaps even overclose the universe.
For this reason, let us estimate how long it would take for percolation to occur, assuming the existence of a network from the time $t_{0}$. For a fixed final time $t$, the length distribution of the loop population can be derived by the production rate (3.19), as well as the fact that a loop produced at the time $t_{p}$ will have grown as
$\ell(t)=\alpha t_{p}\left(\frac{t}{t_{p}}\right)^{m/2},$ (4.34)
@@
$\frac{{\rm d}n}{{\rm d}\ell}=\frac{2\mathcal{F}_{\alpha}\alpha^{\frac{6}{2-m}}}{(2-m)}\frac{c\bar{v}}{\gamma\xi^{3}}\frac{t^{\frac{3m}{2-m}}}{\ell^{\frac{8-m}{2-m}}}.$ (4.35)
As in [35], percolation may be estimated to have occured when the typical string separation is the same order as the string length, namely
$\int_{\ell(t_{0})}^{\ell(t)}{{\rm d}\ell}\,\frac{\ell(t)^{3}}{a(t)^{3}}\frac{{\rm d}n}{{\rm d}\ell}\sim\mathcal{O}(1).$ (4.36)
@@
$\frac{t}{t_{0}}\lesssim\left[\frac{3m}{2\mathcal{F}_{\alpha}}\frac{\gamma\xi^{3}}{c\bar{v}}\alpha^{\frac{3m-6}{2-m}}\right]^{\frac{2n}{3(mn-4)}}.$ (4.37)
@@
$\frac{t}{t_{0}}\lesssim 10^{6}-10^{7}.$ (4.38)
This is a sizeable fraction (on a log scale) of the maximal allowed lenght of kination in LVS, $\frac{t}{t_{0}}\sim 10^{10}$ [35].
The question of what happens afterwards is slightly more subtle, and we suspect it cannot be answered with the tools at our disposal. [^22] Progress on this matter could be achieved using the Boltzmann equation approach of [100, 56]. Therefore, this paragraph should be considered mainly as speculative. As the loops merge with one another and start to percolate, they start to feed back into the population of long strings. To model this within the VOS framework, a cheap attempt may be to decrease the loop chopping efficiency $\tilde{c}$ governing the decays of long strings into loops. If $\tilde{c}$ remains lower bounded, then the network will also be exhibiting a quasi-scaling regime as per Eq. (D.2). On the other hand, $L(t)$ may grow sub-linearly if $\tilde{c}\rightarrow 0$. Even if this is the case, the network can never dominate the energy density of the universe, unlike what would happen with a constant tension. Indeed, by (D.2),
$\rho=\frac{\mu}{L^{2}}\leq\frac{C}{t^{2}},$ (4.39)